There is a 4 foot high fence that surrounds your backyard. The fence casts a shadow that is 5 feet long. If you are 6 feet tall and standing next to the fence, how long will your shadow be?
Answer by TimothyLamb(4379)

s/h = shadow/height
s/h = 5/4
---
s/6 = 5/4
s = 6*5/4
6*5/4
---
copy and paste the above expression in to this calculator:
http://sooeet.com/math/scientific-calculator.php
---
your shadow = 7.5 feet
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---